The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

         mcnews@astro.uiuc.edu
http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
 http://www.astro.uni-bonn.de/~mcnews/

No. 36 September 15, 1999

Contents

News  1
Abstracts of 9 refereed papers 2
3 conference contributions  8
Job opportunity  10

News

Catalog of Cepheids from the LMC available over the Internet

The OGLE-II microlensing search team announces release of the OGLE Catalog of Cepheids from the Large Magellanic Cloud. The catalog contains 1333 objects and more than 3.4 · 10^{5} BVI measurements of these stars.

The Catalog, finding charts and BVI individual observations for all objects are available from the OGLE Internet archive at

http://www.astrouw.edu.pl/~ftp/ogle

or its US mirror at

http://www.astro.princeton.edu/~ogle

Andrzej Udalski

Abstracts of Refereed Papers

Dust and Stellar Populations in the Large Magellanic Cloud

Dennis Zaritsky^{1,2}

1 UCO/Lick Observatory, Santa Cruz, CA, USA 2 Steward Observatory, Tucson, AZ, USA

We present an analysis of line-of-sight extinction measurements obtained using data from the Magellanic Clouds Photometric Survey (Zaritsky, Harris, & Thompson 1997, AJ, 114, 1002), which provides 4-filter photometry for millions of stars in the Large Magellanic Cloud. We find that visual extinctions are typically larger by several tenths of a magnitude for stars with effective temperatures > 12000 K, than for stars with effective temperatures between 5500 K and 6500 K. Several repercussions of this population-dependent extinction are discussed. In particular, LMC distance measurements that utilize old stellar populations, but use extinctions derived from OB stars, may be biased low. As a specific example, we show that the LMC distance modulus derived from field red clump stars is revised upward relative to published measurements by ∼ 0.2 mag if one uses the extinction measured for a matched stellar population. Conversely, measurements that utilize the youngest stars are subject to greater, and more variable, extinction leading preferentially to results that may be biased high. Population-dependent extinction affects the interpretation of color-magnitude diagrams and results in an effective absorption law that is steeper than that intrinsic to the dust for unresolved stellar systems. We further explore the relation between the stellar populations and dust by comparing our extinction map to the 100μm image of the region and identifying potential heating sources of the dust. We find that although regions of high 100μm flux are associated with young stars, young stars are not necessarily associated with regions of high 100μm flux and that ∼ 50% of the 100μm flux is emitted beyond the immediate regions of high OB stellar density. We conclude that 100μm flux should be used with caution as a star formation tracer, particularly for studies of star formation within galaxies. Finally, we reproduce the observed extinction variation between the hot and cold stellar populations with a simple model of the distribution of the stars and dust where the scaleheight of the cooler stars is ≫ than that of the dust (which is twice that of the OB stars; Harris, Zaritsky, & Thompson 1997, AJ, 114, 1933).

Accepted by: the Astronomical Journal

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9908363

ORFEUS II echelle spectra: on the H_{2}/CO ratio in the LMC towards LH 10

P. Richter^{1}, K.S. de Boer^{1}, D.J. Bomans^{2}, Y.-N. Chin^{3,4}, A. Heithausen^{4}, J. Koornneef^{5}

1 Sternwarte, Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 2 Astronomisches Institut, Ruhr-Universität Bochum, Postfach 102148, D-44780 Bochum, Germany 3 Institute of Astronomy and Astrophysics, Academia Sinica, P.O.Box 1-87, Nankang, 11529 Taipei, Taiwan 4 Radioastronomisches Institut, Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 5 Kapteyn Institute, Postbus 800, NL-9700AV Groningen, The Netherlands

We have investigated the H_{2}/CO abundance ratio in diffuse LMC gas in direction of the star 3120 in the association LH 10. ORFEUS II echelle spectra were used to determine N(H_{2}), N(H i) and the upper limit of N(CO) from the analysis of their absorption lines in the FUV domain. We find N(H_{2})/N(^{12}CO) ≥ 2.0 × 10^{5} as a lower limit for the crucial H_{2}/CO ratio. Additional SEST measurements of CO emission in that direction show that the bulk of CO must be located behind the target star LH 10:3120. The molecular fraction in the LMC gas is low and we speculate that it is limited by the lower dust content in the LMC.

Accepted by: Astronomy & Astrophysics

For preprints, contact prichter@astro.uni-bonn.de

Gamma Ray Bursts versus OB Associations: Do they trigger star formation?

Yu. N. Efremov^{1}, S. Ehlerová^{2} and J. Palouš^{2}

1 Sternberg Astronomical Institute, MSU, Moscow 119899, Russia 2 Astronomical Institute, Academy of Sciences of the Czech Republic, Boční II 1401, 141 31 Prague 4, Czech Republic

We discuss differences in shapes, expansion velocities and fragmentation times of structures created by an energy deposition from a single Gamma Ray Burst and an OB association to the ISM. After the initial inflation, supershells produced by GRBs are almost static or slowly expanding, contrary to more rapidly expanding supershells created by OB associations. We discuss the position of the energy source relative to the symmetry plane of the galaxy: observed arc-like structures can be the most dense part of structures formed by an expansion from a source above or below the galactic plane. Arcs may also form if the expansion takes place inside a giant H i cloud. We try to reproduce the size, the age, and the average distance between OB associations in the Sextant region at the edge of LMC 4.

Accepted by: Astronomy & Astrophysics For preprints, contact sona@ig.cas.cz

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9908044

The LMC eclipsing binary HV 2274: fundamental properties and comparison with evolutionary models

I. Ribas^{1}, E.F. Guinan^{2}, E.L. Fitzpatrick^{2}, L.E. DeWarf^{2}, F.P. Maloney^{2}, P.A. Maurone^{2}, D.H. Bradstreet^{3}, A. Giménez^{4}, and J.D. Pritchard^{5}

1 Departament d’Astronomia i Meteorologia, Universitat de Barcelona, Av. Diagonal, 647, E-08028 Barcelona, Spain 2 Department of Astronomy and Astrophysics, Villanova University, Villanova, PA 19085, USA 3 Department of Physical Science, Eastern College, St. Davids, PA 19087, USA 4 LAEFF (INTA-CSIC), Apartado 50727, E-28080 Madrid, Spain 5 Department of Physics & Astronomy, University of Canterbury, Private Bag 4800, Christchurch, New Zealand

We are carrying out an international, multi-wavelength program to determine the fundamental properties and independent distance estimates of selected 14^{th} to 15^{th} magnitude eclipsing binaries in the Large and Small Magellanic Clouds (LMC and SMC). Eclipsing binaries with well-defined double-line radial velocity curves and light curves provide valuable information on orbital and physical properties of their component stars. These properties include, among other characteristics, stellar mass and radius. These can be measured with an accuracy and directness unachievable by any other means. The study of stars in the LMC and SMC where the metal abundances are significantly lower than solar (by 1/3 to 1/10) provides an important opportunity to test stellar atmosphere, interior and evolution models, and opacities. For the first time, we can also measure direct Mass-Luminosity relations for stars outside our Galaxy.

In a previous paper we demonstrated how a precise distance to the LMC — corresponding to (V_{∘} − M_{v})_{LMC} = 18.30 ± 0.07 mag — could be determined using the 14^{th} mag LMC eclipsing binary HV 2274. In this paper we concentrate on the determination of the orbital and physical properties of HV 2274 and its component stars from analyses of light curves and new radial velocity curves formed from HST/Goddard High-Resolution Spectrograph observations. HV 2274 (B1-2 IV-III + B1-2 IV-III; V_{max} ≈ +14.2; P = 5^{d}.73) is a particularly appealing star because it is a detached binary which has an eccentric orbit (e = 0.136) and shows fast apsidal motion. The results of these analyses yield reliable masses and absolute radii, as well as other physical and orbital properties of the stars and the system. From UV/optical spectrophotometry (1150−4820 Å) of HV 2274 obtained with the HST/Faint Object Spectrograph, the temperatures and the metallicity ([F e/H] = −0.45 ± 0.06) of the stars were found, as well as the interstellar extinction of the system. The values of mass, absolute radius, and effective temperature, for the primary and secondary stars are: 12.2 ± 0.7 M_{⊙}, 9.9 ± 0.2 R_{⊙}, 23000 ± 180 K, and 11.4 ± 0.7 M_{⊙}, 9.0 ± 0.2 R_{⊙}, 23110 ± 180 K, respectively.

The age of the system (τ = 17 ± 2 Myr), helium abundance (Y = 0.26 ± 0.03) and a lower limit of the convective core overshooting parameter of α_{ov} ≈ 0.2 were obtained from fitting the stellar data with evolution models of Claret & Giménez. HV 2274 has a relatively well-determined (and fast) apsidal motion period of U = 123 ± 3 yr. From the analysis of apsidal motion, additional information and constraints on the structure of the stars can be made. The apsidal motion analysis corroborates that some amount of convective core overshooting (α_{ov} between 0.2 and 0.5) is needed.

Accepted by: The Astrophysical Journal For preprints, contact iribas@am.ub.es

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9909046

Early-Type Contact Systems in the LMC MACHO Database

S. M. Rucinski^{1}

1 David Dunlap Observatory, University of Toronto, P.O.Box 360, Richmond Hill, Ontario L4C 4Y6, Canada

Two sub-samples of blue, luminous contact systems from among 86 systems discovered in LMC by the MACHO project and classified as EB3 have been analyzed: a sub-sample of 29 extremely-blue systems (XB) with the observed colors V − R_{C} < −0.1 and a sub-sample of 36 moderately-blue (MB) systems with −0.1 < V − R_{C} < 0. To be so blue, the XB systems must be intrinsically very hot and almost unreddened so that lack of information on reddening is unimportant for them; their properties offer us a first-time insight into the absolute-magnitude calibration for massive contact binaries. It has been found that the LMC systems are apparently not limited by the blue, short-period envelope in the color-period diagram which had been established on the basis of systems observed in the solar neighborhood, in galactic open clusters and in the sample of Old Disk systems in the direction of the Galactic Bulge. The period–magnitude correlation for the XB systems has a similar slope to those established in the absolute-magnitude, B − V and V − I_{C} based, calibrations for W UMa-type systems.

Accepted by: Acta Astronomica

For preprints, contact rucinski@astro.utoronto.ca
Also available from the URL http://www.astro.utoronto.ca/∼rucinski/rucinski.html

Optical spectroscopy of the candidate luminous white dwarf in the young LMC cluster NGC 1818

M.R. Burleigh^{1}, R.A. Saffer^{2}, G.F. Gilmore^{3}, R. Napiwotzki^{4}

1 Dept. of Physics and Astronomy, University of Leicester, Leicester LE1 7RH, UK 2 Dept. of Astronomy and Astrophysics, Villanova University, PA 19085, USA 3 Institute of Astronomy, Madingley Road, Cambridge, CB3 0HA, UK 4 Dr. Remeis-Sternwarte, Sternwartstr. 7, D-96049 Bamberg, Germany

An optical spectrum of the Elson et al. (1998) candidate luminous white dwarf in the young LMC cluster NGC 1818 shows conclusively that it is not a degenerate star. A model atmosphere fit gives T≈31,500K and log g≈4.4, typical of a garden-variety main sequence B star. However, if it is a true LMC member then the star is under-luminous by almost three magnitudes. Its position in the cluster colour-magnitude diagram also rules out the possibility that this is an ordinary B star. The luminosity is, however, consistent with a 0.5M_{⊙} post-AGB or post-EHB object, although if it has evolved via single star evolution from a high mass (7.6−9.0M_{⊙}) progenitor then we might expect it to have a much higher mass, ∼0.9M_{⊙}. Alternatively, it has evolved in a close binary. In this case the object offers no implications for the maximum mass for white dwarf progenitors, or the initial-final mass relation. Finally, we suggest that it could in fact be an evolved member of the LMC disk, and merely projected by chance onto NGC 1818. Spectroscopically, though, we cannot distinguish between these evolutionary states without higher resolution (echelle) data.

Accepted by: Monthly Notices of the Royal Astronomical Society (pink pages) For preprints, contact mbu@star.le.ac.uk

Also available from the URL http://ledas-www.star.le.ac.uk/Preprint/

The Optical Gravitational Lensing Experiment. Cepheids in the Magellanic Clouds. IV. Catalog of Cepheids from the Large Magellanic Cloud

A. Udalski^{1}, I. Soszyński^{1}, M. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1}, P. Woźniak^{2}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present the Catalog of Cepheids from the LMC. The Catalog contains 1333 objects detected in the 4.5 square degree area of central parts of the LMC. About 3.4 · 10^{5} BVI measurements of these stars were collected during the OGLE-II microlensing survey. The Catalog data include period, BVI photometry, astrometry, and R_{21}, ϕ_{21} parameters of the Fourier decomposition of I-band light curve.

The vast majority of objects from the Catalog are the classical Cepheids pulsating in the fundamental or first overtone mode. The remaining objects include Population II Cepheids and red giants with pulsation-like light curves.

Tests of completeness performed in overlapping parts of adjacent fields indicate that completeness of the Catalog is very high: > 96%. Statistics and distributions of basic parameters of Cepheids are also presented.

Finally, we show the light curves of three eclipsing systems containing Cepheid detected among objects of the Catalog.

All presented data, including individual BVI observations are available from the OGLE Internet archive.

Submitted to: Acta Astronomica Vol. 49

For preprints, contact udalski@sirius.astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9908317

UBV I Color-Magnitude Diagrams in Baade’s Window:

Metallicity Range and the Implications for the Red Clump Method and the Distance to the Large Magellanic Cloud

K.Z. Stanek^{1}, J. Kaluzny^{2}, A. Wysocka^{2}, I. Thompson^{3}

1 Harvard-Smithsonian Center for Astrophysics, USA 2 Warsaw University Observatory, Poland 3 Carnegie Observatories, USA

We analyze the UBV I color-magnitude diagrams of a relatively low-reddening region of Baade’s Window. The dereddened red giant branch is very wide [∼ 1.0 mag in (U − B)_{0} and ∼ 0.4 mag in (B −V)_{0} and (V −I)_{0}], indicating a significant dispersion of stellar metallicities, which by comparison with the theoretical isochrones and data for Galactic clusters we estimate to lie between approximately −0.8 < [F e/H] < +0.3, i.e. spanning at least 1 dex in metallicity, in good agreement with earlier spectroscopic studies.

We use these data to discuss the metallicity dependence of the red clump I-band brightness and we show that it is about 0.1 mag dex^{−1}, in very good agreement with the previous empirical determinations and the models of stellar evolution.

We also discuss the red clump distance to the Large Magellanic Cloud and argue that the value of distance modulus μ_{0,LMC} = 18.18 ± 0.06, determined by Udalski, is currently the most secure and robust of all LMC distance estimates. This has the effect of increasing any LMC-tied Hubble constant by 15%, including not only the recent determination by the HST Key Project, but also that of Sandage et al.

Submitted to: The Astronomical Journal

For preprints, contact kstanek@cfa.harvard.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9908041

Self-Lensing Models of the LMC

G. Gyuk^{1}, N. Dalal^{1} and K. Griest^{1}

1 Physics Dept., UCSD, La Jolla CA 92093, USA

All of the proposed explanations for the microlensing events observed towards the LMC have difficulties. One of these proposed explanations, LMC self-lensing, which invokes ordinary LMC stars as the long sought-after lenses, has recently gained considerable popularity as a possible solution to the microlensing conundrum. In this paper, we carefully examine the set of LMC self-lensing models. In particular, we review the pertinent observations made of the LMC, and show how these observations place limits on such self-lensing models. We find that, given current observational constraints, no purely LMC disk models are capable of producing optical depths as large as that reported in the MACHO collaboration 2-year analysis. Besides pure disk, we also consider alternate geometries, and present a framework which encompasses the previous studies of LMC self-lensing. We discuss which model parameters need to be pushed in order for such models to succeed. For example, like previous workers, we find that an LMC halo geometry may be able to explain the observed events. However, since all known LMC tracer stellar populations exhibit disk-like kinematics, such models will have difficulty being reconciled with observations. For SMC self-lensing, we find predicted optical depths differing from previous results, but more than sufficient to explain all observed SMC microlensing. In contrast, for the LMC we find a self-lensing optical depth contribution between 0.47 · 10^{−8} and 7.84 · 10^{−8}, with 2.44 · 10^{−8} being the value for the set of LMC parameters most consistent with current observations.

Submitted to: The Astrophysical Journal

For preprints, contact endall@physics.ucsd.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/?9907338

Abstracts of Non-Refereed Papers

The Magellanic Stream: RATAN-600 Studies According to an Extended Program

N.V. Bystrova ^{1}

1 St. Petersburg Branch of the Special Astrophysical Observatory, Russian Ac. Sci. Pulkovo, 196140, St. Petersburg, Russia

According to the data from the Pulkovo H i Sky Survey around the Magellanic Stream’s great circle the extended areas of H i emission were possibly found at rather small negative LSR velocities. In order to check the existence of the above neutral hydrogen regions special observations at the RATAN-600 telescope were made.

The radiotelescope RATAN-600 situated in Zelenchukskaja, Karachai-Cherkessia, has a fan beam of 2.5 arcmin × 2.5 deg. Its 40 spectral channels cover the velocity range of about 250 km/s, the bandwidth and spacing of the channels are 30 kHz. The series of observations were as 6 or 7 hours long drift scans north from declination −40 deg in 5 deg intervals and on several special levels. During the 4.5 years (1995–99) 9 sessions of observations were realized and the reductions for 7 of them are already made. Apart from details of the structural component of the neutral hydrogen emission broad signals with halfwidths of more than two hours and amplitude of several K-s are also seen. These broad emissions are seen at least in 5 – 6 channels in the RATAN-600 cuts across the Magellanic Stream zone. This yields the velocity dispersion for the broad emission areas as 30 – 40 km/s. As seen in the records these broad signals displace themselves for more negative velocities in the RA direction as does the galactic emission forming the N-warp of the neutral hydrogen gas layer. But the displacement in the case of Magellanic Stream is several times more than is seen in the Galaxy.

We have taken the azimuthally opposite region on the great circle of the Magellanic Stream in order to have for registration the descent part of the warp (if any). The detailed discussion is under way now but such preliminary statements could be made: There is the neutral hydrogen emission on the Magellanic Stream’s great circle at least in the second region of the Stream studied. The type of signals is almost the same as in the first region – their velocities are also rather negative. But no displacement of the emission in the RA direction for more negative velocity was found here. Either no displacement is seen at all or it is in opposite direction (opposite to RA direction). The scale of displacements is less than in the first region of the Stream studied. After the reductions of the full material of observations including the data obtained in 1999 more accurate conclusions could be possibly made.

To appear in: Conference proceedings of the Gamov Memorial International Conference “Early Universe: Cosmological Problems and Instrumental Technologies”, St. Petersburg, August 23 – 27, 1999, in the Bull. of the Special Astrophys. Obs. For preprints, contact bnv@fsao.spb.su

High-Velocity Clouds of the Magellanic System

M.E. Putman^{1}

1 Research School of Astronomy & Astrophysics, Australian National University, Weston Creek P.O., Weston, ACT 2611, Australia

The results of the interaction between the Milky Way and the Magellanic Clouds are revealed through several high velocity complexes which are connected to the Clouds. The exact mechanism of their formation is under some debate, but they remain the only group of high-velocity clouds (HVCs) for which we have an origin (and roughly a distance). Given that, the Magellanic HVCs can be used as a calibrator for other HVCs, while also providing an opportunity to closely investigate the remnants of an interacting system. These HVCs may hold the key to the star formation history, kinematic structure and present Hubble type of the Magellanic Clouds, and their proximity to the Milky Way allows us to estimate key Galactic parameters.

The HVCs of the Magellanic System can be classified into three major complexes: the Magellanic Stream, the largest H i feature in the sky outside of our Galaxy which trails the Clouds; the Magellanic Bridge, an H i connection between the Clouds; and the Leading Arm, a more diffuse H i filament which leads the Clouds. In terms of HVCs, these features have been studied rather extensively. In this review, I will first describe the observational results for each complex and subsequently discuss their origin and relationship to the overall HVC population. All of the H i data presented is from the H i Parkes All-Sky Survey (HIPASS) (Staveley-Smith 1997; Putman & Gibson 1999).

To appear in: “High-Velocity Clouds”, contributed chapter, (Kluwer: Dordrecht), edited by H. van Woerden, U. Schwarz, B. Wakker and K.S. de Boer

For preprints, contact putman@mso.anu.edu.au
Also available from the URL http://msowww.anu.edu.au/~putman/pubs.html

HST study of the Stellar Populations within 30 pc of SN 1987A

M. Romaniello^{1}, N. Panagia^{2}, and S. Scuderi^{3}

1 ESO, Karl-Schwarzschild-Strasse 2, D-85748 Garching bei München, Germany. 2 STScI, 3700 San Martin Drive, Baltimore, MD 21218, USA. Affiliated to the Astrophysics Division, Space Science Department of ESA. 3 Osservatorio di Catania, Viale A. Doria 6, Catania, Italy.

We have studied the stellar populations in a region of 30 pc around SN 1987A in the Large Magellanic Cloud using multi-band HST -WFPC2 images.

The effective temperature, luminosity and reddening of each detected star were determined by fitting the measured broad band magnitudes to the ones calculated with model atmospheres. The resulting HR diagram reveals the presence of stars with ages between 1 and 150 Myrs, superposed on a much older field population. The youngest stars in the field appear to be T Tauri stars, characterized by strong Hα excesses.

The star formation rate has increased monotonically in the last 8 Gyrs and the star formation activity is still very high at present.

Young stars of low and high mass have different spatial distributions. The latter ones are strongly concentrated in the vicinities of the Supernova, whereas the previous ones are more evenly distributed on the field. Hence, the Mass Function varies on a scale of a few parsecs. However, averaging over the entire area, one can define an Initial Mass Function, well fitted by a power-law with a slope α < 2.55. The uncertainty on the IMF due to incompleteness in identifying T Tauri stars is discussed.

To appear in: “Star Formation 1999”, Nagoya University (Japan), 21-25 June 1999. For preprints, contact mromanie@eso.org

Job Opportunity

Postdoctoral Research Position at the Max Planck Institute for Astronomy, Heidelberg

The Max-Planck-Institut für Astronomie (MPIA) in Heidelberg, Germany, invites applications for a postdoctoral research position in the general area of stellar population studies and dwarf galaxies with Dr. Eva Grebel. Possible research topics include population synthesis modelling, analysis of wide-field surveys of individual galaxies or nearby groups (including SDSS data), abundance studies, or kinematic studies based on optical/H i data. Experience with optical/infrared observations and analysis techniques will be an asset. Experience with other wavelength regions is welcome.

The MPIA operates Calar Alto Observatory (http://www.mpia-hd.mpg.de/CAHA/), has access to ESO (La Silla and VLT) and time at UKIRT, participates in the SDSS Project, and has GRAPE hardware and supercomputing facilities. More about the MPIA and current staff interests can be

found at the URL http://www.mpia-hd.mpg.de/welcome.html.

Applicants should have a Ph.D. in astronomy, physics, or a related field by the start of the appointment and assume the position in the fall of 2000; earlier or later starting dates may be negotiated. The position is offered initially for two years with possible extensions up to five years. The salary will be on the German BAT II scale and depends on age and marital status.

Applicants should send a curriculum vitae, bibliography, statement of research interests and experience, and arrange for three letters of recommendation to be sent to

Dr. Eva K. Grebel

Max-Planck-Institut für Astronomie

Königsstuhl 17

D–69117 Heidelberg

Germany

by December 1, 1999. Questions may be addressed to Eva Grebel by e-mail (grebel@mpia-hd.mpg.de or grebel@astro.washington.edu) or during the winter AAS meeting.